Find the complete integral of the PDE px+qy+z=xq²

Explain the problem with step by step process?


1
Expert's answer
2021-09-25T11:46:13-0400

The provided equation is: PDE "px+qy+z=xq\u00b2"

We can use Charpit's auxiliary equations to get: 

"ds= \frac{dp}{0}= \frac{dq}{0}=\frac{dz}{z+pq}=\frac{dx}{x+q} =\frac{dy}{y+p}\\

ds=\frac{dp}{0}, \frac{dp}{0}=\frac{dq}{0} \implies p=C, q=D"

We get a complete integral of: 

"dz = pdx + qdy = Cdx + Ddy \\

z(x,y) = Cx + Dy + E"